# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v2

@@ B.2 Lower bound $\widetilde{\alpha}_{0}(\mathbf{x})$
We proceed by taking expectation of the right-hand side of (E.13) with respect to $\tilde{\xi}_{1},\dotsc,\tilde{\xi}_{n}$. Define the shorthand $\lambda:=\frac{\gamma^{4}}{4n\lambda_{\max}^{2}\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}$. Fix an arbitrary $\lambda_{0}>0$ for now (we will determine its value later) and define truncated parameter $\bar{\lambda}=\min(\lambda,\lambda_{0})$. We have
$\operatorname{\mathbb{E}}\left[e^{-\lambda(\sum_{i}\tilde{\xi}_{i})^{2}}\right]$ $\leq\operatorname{\mathbb{E}}\left[e^{-\bar{\lambda}(\sum_{i}\tilde{\xi}_{i})^{2}}\right]$ (E.14)
@@
$=e^{-\bar{\lambda}n^{2}}\left(1-p+pe^{2\bar{\lambda}+\bar{\lambda}^{2}/2}\right)^{n(n-1)}\,.$ (E.15)
@@
We next further upper bound the right-hand side of E.15 to get a simpler expression. In doing that we use the following lemma.
Lemma E.3
For $t\leq 1/2$ we have $e^{t}-1\leq t+t^{2}$.
By choosing $\lambda_{0}\leq 0.2$ we have $22\lambda_{0}+\lambda_{0}^{2}/2\leq 1/$ 2 and so $22\bar{\lambda}+\bar{\lambda}^{2}/2\leq 1/$ 2. By virtue of the above lemma we have
@@
By using this bound in (E.15) we obtain
$\operatorname{\mathbb{E}}\left[e^{-\lambda(\sum_{i}\tilde{\xi}_{i})^{2}}\right]\leq e^{-\bar{\lambda}n^{2}}e^{(2+6\lambda_{0})\bar{\lambda}pn^{2}}=e^{-(1-2p^{\prime})\bar{\lambda}n^{2}}\,,$ (E.16)
@@
Using the above in (E.13) and then taking expectation with respect to ${\bm{z}}_{\ell}$ we get
@@
$\leq\frac{1}{2}\operatorname{\mathbb{E}}\left[\exp\left(-2(1-2p^{\prime})\bar{\lambda}n^{2}\right)\right]^{1/2}\operatorname{\mathbb{E}}\left[\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{n}\right)\right]^{1/2}\,,$ (E.17)
@@
$\operatorname{\mathbb{E}}\left[\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{n}\right)\right]=\left(1-\frac{2}{n}\right)^{-d/2}\leq e^{2d/n}\,,$ (E.18)
@@
We continue by bounding the first term on the right-hand side of (E.17). Define the event $\mathcal{E}_{1}$ as follows:
$\mathcal{E}_{1}:=\left\{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}\leq\sqrt{2d}\right\}\,.$ (E.19)
@@
$\leq\exp\left(-\frac{\gamma^{4}}{4\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+\mathbb{P}(\mathcal{E}_{1}^{c})$ (E.20)
Note that in step $(a)$, we used the fact that on the event $\mathcal{E}_{1}$, we have $\lambda\geq{\gamma^{4}}/(8\lambda_{\max}^{2}nd)$ and we choose $\lambda_{0}={\gamma^{4}}/(8\lambda_{\max}^{2}nd)$. This way, we have $\bar{\lambda}=\min(\lambda,\lambda_{0})={\gamma^{4}}/(8\lambda_{\max}^{2}nd)$. Also note that by our assumption in the statement of the lemma, we have $\lambda_{0}\leq 0.2$ which is the condition assumed in deriving (E.16).
We next bound $\mathbb{P}(\mathcal{E}_{1}^{c})$. Since ${\bm{z}}_{\ell}\sim{\mathcal{N}}(0,\bm{I}_{d})$, $\left\|\bm{z}_{\ell}\right\|_{\ell_{2}}^{2}$ is $\chi^{2}$ distribution with $d$-degrees of freedom. Using the tail bound of $\chi^{2}$ distribution (see e.g, (Ghosh,, 2021, Theorem 3)) we have $\mathbb{P}(\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}\geq 2d)\leq e^{-d/8}$ and so
@@
Using the above in (E.20) we obtain
$\operatorname{\mathbb{E}}\left[\exp\left(-2(1-2p^{\prime})\bar{\lambda}n^{2}\right)\right]$ $\leq\exp\left(-\frac{\gamma^{4}}{4\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/8}.$ (E.21)
By combining (E.21) and (E.18) with (E.17) we arrive at
@@
E.2 Proof of Lemma E.2
Under our data model we have $\bm{x}_{\ell}=y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}$. Substituting for $\bm{x}_{\ell}$ in the expression of $T_{\ell}$, we have
@@
$\operatorname{\mathbb{E}}[T_{\ell}|\beta]=(1-2q_{0})\langle y\bm{x},(1+u_{\ell}){\bm{{\mu}}}\rangle\geq(1-2q_{0})\langle y\bm{x},{\bm{{\mu}}}\rangle\,,$
@@
if $\frac{\beta\gamma^{2}}{2d\lambda_{\max}}>1$, where the last step follows from the observation that $\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}\sim\chi^{2}_{d}$ and using the tail bound of $\chi^{2}_{d}$(Ghosh,, 2021, Theorem 3)). So by definition of $q_{0}$ we showed that
$q_{0}\leq q:=\exp\left(-\frac{\beta\gamma^{2}}{20\lambda_{\max}}\right)\leq e^{-d/10}<\frac{1}{2}\,,$
which completes the proof of lemma.
Appendix F Proof of Theorem 4.9
Using Theorem 4.8, we have ${\sf acc}({\widehat{\bm{\theta}}}_{1})\geq\operatorname{\mathbb{E}}[\alpha_{1}({\bm{x}})]$. Based on our data model, $\bm{x}=y(1+u){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}$. Consider the probabilistic event $\mathcal{E}_{2}:=\{{\bm{z}}:\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\leq 2d\}$. For every $\eta>0$, we have
@@
$\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}=\left\|\bm{\Sigma}{\bm{z}}\right\|_{\ell_{2}}^{2}\leq\lambda_{\max}^{2}\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\leq 2\lambda_{\max}^{2}d$
under the event $\mathcal{E}_{2}$. Further, $\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\sim\chi^{2}_{d}$ and so using tail bounds for $\chi^{2}_{d}$, see e.g., (Ghosh,, 2021, Theorem 3), we have $\mathbb{P}(\mathcal{E}_{2}^{c})\leq e^{-d/8}$. Putting things together, we obtain
@@
Appendix G Remaining Experimental Details
Here we provide the remaining details about the experiments in Section 5. Our experiments were done using TensorFlow and JAX on one 40 GB A100 GPU (per run). In all the cases, we retrain starting from random initialization rather than the previous checkpoint we converged to before RT; the former worked better than the latter. We list training details for each individual dataset next.
CIFAR-10. Optimizer is SGD with momentum $=0.9$, batch-size $=32$, number of gradient steps in each stage of training (i.e., both stages of baseline, full RT and consensus-based RT) $=21$k. We use the cosine one-cycle learning rate schedule with initial learning rate $=0.1$ for each stage of training. The number of gradient steps and initial learning rate were chosen based on the performance of the baseline method and not based on the performance of full or consensus-based RT. Standard augmentations such as random cropping, flipping and brightness/contrast change were used.
CIFAR-100. Details are the same as CIFAR-10 except that here the number of gradient steps in each stage of training $=28$k and initial learning rate $=0.005$.
@@ Appendix G Remaining Experimental Details
Appendix H Consensus-Based Retraining Does Better than Confidence-Based Retraining
Here we compare full and consensus-based RT against another strategy for retraining which we call confidence-based retraining (RT). Specifically, we propose to retrain with the predicted labels of the samples with the top $50\%$ margin (i.e., highest predicted probability - second highest predicted probability); margin is a measure of the model’s confidence. This idea is similar to self-training’s method of sample selection in the semi-supervised setting (Amini et al.,, 2022). In Tables 8 and 9, we show results for CIFAR-10 and CIFAR-100 (in the same setting as Section 5 and Appendix G) with the smallest value of $\epsilon$ from Tables 1 and 2, respectively. Notice that consensus-based RT is clearly better than confidence-based RT.
@@ Appendix H Consensus-Based Retraining Does Better than Confidence-Based Retraining
Appendix I Retraining is Beneficial Even Without a Validation Set
In all our previous experiments, we assumed access to a small clean validation set. Here we consider training ResNet-34 on CIFAR-100 when we do not have access to a validation set and keep training for a fixed number of epochs (100). The point of these experiments is to show that retraining can offer gains even without a validation set, in which case we can expect severe overfitting. Please see the the results and discussion in Table 10.
Table 10: No validation set (CIFAR-100 w/ ResNet-34). Test set accuracies (mean $\pm$ standard deviation). Just like all our previous results in the presence of a small validation set, consensus-based RT is the clear winner here. However, due to a higher degree of overfitting here, the performance and amount of improvement obtained with retraining is worse than the corresponding experiment using a validation set (Table 4). This is not surprising.
$\epsilon$ Baseline Full RT Consensus-based RT
3 $14.53\pm 0.48$ $15.20\pm 0.71$ $\textbf{17.30}\pm 0.43$
@@ Appendix I Retraining is Beneficial Even Without a Validation Set
Appendix J Beyond Label DP: Evaluating Retraining in the Presence of Human Annotation Errors
Even though our empirical focus in this paper has been label DP training, retraining (RT) can be employed for general problems with label noise. Here we evaluate RT in a setting with “real” label noise due to human annotation. Specifically, we focus on training a ResNet-18 model (without label DP to be clear) on the CIFAR-100N dataset introduced by Wei et al., (2021) and available on the TensorFlow website. CIFAR-100N is just CIFAR-100 labeled by humans; thus, it has real human annotation errors. The experimental setup and details are the same as CIFAR-100 (as stated in Section 5 and Appendix G); the only difference is that here we use initial learning rate $=0.01$.
In Table 11, we list the test accuracies of the baseline which is just vanilla training with the given labels, full RT and consensus-based RT, respectively. Even here with human annotation errors, consensus-based RT is beneficial.

### v3

@@ Appendix E Comparison with Other Lower Bounds for Learning with Noisy Labels
We proceed by taking expectation of the right-hand side of (F.14) with respect to $\tilde{\xi}_{1},\dotsc,\tilde{\xi}_{n}$. Define the shorthand $\lambda:=\frac{\gamma^{4}}{4n\lambda_{\max}^{2}\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}$. Fix an arbitrary $\lambda_{0}>0$ for now (we will determine its value later) and define truncated parameter $\bar{\lambda}=\min(\lambda,\lambda_{0})$. We have
$\operatorname{\mathbb{E}}\left[e^{-\lambda(\sum_{i}\tilde{\xi}_{i})^{2}}\right]$ $\leq\operatorname{\mathbb{E}}\left[e^{-\bar{\lambda}(\sum_{i}\tilde{\xi}_{i})^{2}}\right]$ (F.15)
@@
$=e^{-\bar{\lambda}n^{2}}\left(1-p+pe^{2\bar{\lambda}+\bar{\lambda}^{2}/2}\right)^{n(n-1)}\,.$ (F.16)
@@
We next further upper bound the right-hand side of F.16 to get a simpler expression. In doing that we use the following lemma (proved in Section F.1.1).
Lemma F.3.
For $t\leq\frac{1}{2}$, we have $e^{t}-1\leq t+t^{2}$.
By choosing $\lambda_{0}\leq 0.2$ we have $22\lambda_{0}+\lambda_{0}^{2}/2\leq 1/$ 2 and so $22\bar{\lambda}+\bar{\lambda}^{2}/2\leq 1/$ 2. Using the above lemma, we have
@@
By using this bound in (F.16) we obtain
$\operatorname{\mathbb{E}}\left[e^{-\lambda(\sum_{i}\tilde{\xi}_{i})^{2}}\right]\leq e^{-\bar{\lambda}n^{2}}e^{(2+6\lambda_{0})\bar{\lambda}pn^{2}}=e^{-(1-2p^{\prime})\bar{\lambda}n^{2}}\,,$ (F.17)
@@
Using the above in (F.14) and then taking expectation with respect to ${\bm{z}}_{\ell}$ we get
@@
$\leq\frac{1}{2}\operatorname{\mathbb{E}}\left[\exp\left(-2(1-2p^{\prime})\bar{\lambda}n^{2}\right)\right]^{1/2}\operatorname{\mathbb{E}}\left[\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{n}\right)\right]^{1/2}\,,$ (F.18)
@@
$\operatorname{\mathbb{E}}\left[\exp\left(\frac{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}}{n}\right)\right]=\left(1-\frac{2}{n}\right)^{-d/2}\leq e^{2d/n}\,,$ (F.19)
@@
We continue by bounding the first term on the right-hand side of (F.18). Define the event $\mathcal{E}_{1}$ as follows:
$\mathcal{E}_{1}:=\left\{\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}\leq\sqrt{2d}\right\}\,.$ (F.20)
@@
$\leq\exp\left(-\frac{\gamma^{4}}{4\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+\mathbb{P}(\mathcal{E}_{1}^{c})$ (F.21)
Note that in step $(a)$, we used the fact that on the event $\mathcal{E}_{1}$, we have $\lambda\geq{\gamma^{4}}/(8\lambda_{\max}^{2}nd)$ and we choose $\lambda_{0}={\gamma^{4}}/(8\lambda_{\max}^{2}nd)$. This way, we have $\bar{\lambda}=\min(\lambda,\lambda_{0})={\gamma^{4}}/(8\lambda_{\max}^{2}nd)$. Also note that by our assumption in the statement of the lemma, we have $\lambda_{0}\leq 0.2$ which is the condition assumed in deriving (F.17).
We next bound $\mathbb{P}(\mathcal{E}_{1}^{c})$. Since ${\bm{z}}_{\ell}\sim{\mathcal{N}}(0,\bm{I}_{d})$, $\left\|\bm{z}_{\ell}\right\|_{\ell_{2}}^{2}$ has $\chi^{2}$ distribution with $d$-degrees of freedom. Using the tail bound of $\chi^{2}$ distribution (see Theorem 3 of Ghosh (2021)) we have $\mathbb{P}(\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}\geq 2d)\leq e^{-d/8}$ and so
@@
Using the above in (F.21) we obtain
$\operatorname{\mathbb{E}}\left[\exp\left(-2(1-2p^{\prime})\bar{\lambda}n^{2}\right)\right]$ $\leq\exp\left(-\frac{\gamma^{4}}{4\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/8}.$ (F.22)
By combining (F.22) and (F.19) with (F.18) we arrive at
@@
F.1.1 Proof of Lemma F.3
Using the Taylor series expansion of $e^{t}$, we have:
$e^{t}-1$ $=t+\frac{t^{2}}{2}+\sum_{j=3}^{\infty}\frac{t^{j}}{j!}$ (F.23)
$\leq t+\frac{t^{2}}{2}+\frac{t^{2}}{2}\sum_{j=3}^{\infty}\frac{1}{j!}$ (F.24)
$=t+\frac{t^{2}}{2}+\frac{t^{2}}{2}\Big{(}e-\frac{5}{2}\Big{)}$ (F.25)
(F.24) follows by using the fact that for $j\geq 3$ and $t\leq\frac{1}{2}$, $t^{j}\leq t^{3}\leq\frac{t^{2}}{2}$. (F.25) follows by using the fact that $e^{1}-1=1+\frac{1}{2}+\sum_{j=3}^{\infty}\frac{1}{j!}$ after plugging in $t=1$ in Equation F.23.
Next, using the fact that $e<3$ in (F.25), we get:
$e^{t}-1<t+\frac{3t^{2}}{4}<t+t^{2}.$ (F.26)
This completes the proof.
F.2 Proof of Lemma F.2
Under our data model we have $\bm{x}_{\ell}=y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}$. Substituting for $\bm{x}_{\ell}$ in the expression of $T_{\ell}$, we get
@@
$\operatorname{\mathbb{E}}[T_{\ell}|\beta]=(1-2q_{0})\langle y\bm{x},(1+u_{\ell}){\bm{{\mu}}}\rangle\geq(1-2q_{0})\langle y\bm{x},{\bm{{\mu}}}\rangle\,,$ (F.27)
@@
if $\frac{\beta\gamma^{2}}{2d\lambda_{\max}}>1$, where the last step follows from the observation that $\left\|{\bm{z}}_{\ell}\right\|_{\ell_{2}}^{2}\sim\chi^{2}_{d}$ and using the tail bound of $\chi^{2}_{d}$ (Theorem 3 of Ghosh (2021)). So when $\frac{\beta\gamma^{2}}{2d\lambda_{\max}}>1$ and $d\geq 7$,
$q_{0}\leq q:=\exp\left(-\frac{\beta\gamma^{2}}{20\lambda_{\max}}\right)\leq e^{-d/10}<\frac{1}{2}.$
Plugging this into (F.27) completes the proof.
Appendix G Proof of Theorem 4.9
Using Theorem 4.8 (and because the training set $\mathcal{T}$ has measure zero), we have ${\sf err}({\widehat{\bm{\theta}}}_{1})\leq\operatorname{\mathbb{E}}[\alpha_{1}({\bm{x}})]$. Based on our data model, $\bm{x}=y(1+u){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}$. Consider the probabilistic event $\mathcal{E}_{2}:=\{{\bm{z}}:\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\leq 2d\}$. For every $\eta>0$, we have
@@
$\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}=\left\|\bm{\Sigma}{\bm{z}}\right\|_{\ell_{2}}^{2}\leq\lambda_{\max}^{2}\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\leq 2\lambda_{\max}^{2}d,$
under the event $\mathcal{E}_{2}$. Further, $\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\sim\chi^{2}_{d}$ and so using tail bounds for $\chi^{2}_{d}$ (see Theorem 3 of Ghosh (2021)), we have $\mathbb{P}(\mathcal{E}_{2}^{c})\leq e^{-d/8}$. Putting things together, we obtain
@@
Appendix H Remaining Experimental Details
Here we provide the remaining details about the experiments in Section 5. Our experiments were done using TensorFlow and JAX. In all the cases, we retrain starting from random initialization rather than the previous checkpoint we converged to before RT; the former worked better than the latter. We list training details for each individual dataset next.
CIFAR-10. Optimizer is SGD with momentum $=0.9$, batch size $=32$, number of epochs in each stage of training (i.e., both stages of baseline, full RT and consensus-based RT) $=30$. We use the cosine one-cycle learning rate schedule with initial learning rate $=0.1$ for each stage of training. The number of epochs and initial learning rate were chosen based on the performance of the baseline method and not based on the performance of full or consensus-based RT. Standard augmentations such as random cropping, flipping and brightness/contrast change were used.
CIFAR-100. Details are the same as CIFAR-10 except that here the number of epochs in each stage of training $=40$ and initial learning rate $=0.005$.
DomainNet. Details are the same as CIFAR-10 except that here the number of epochs in each stage of training $=15$, initial learning rates for the first stage of the baseline in Tables 7, 8, and 9 are $31e-$ 3, $25e-$ 2, and $45e-$ 4, respectively, whereas the initial learning rate for the second stage of the baseline as well as full RT and consensus-based RT is $31e-$ 3.
@@ Appendix H Remaining Experimental Details
Appendix I Accuracy over the Entire Dataset and over the Consensus Set in the Case of AG News Subset
Here we list the accuracies of the predicted labels and the given labels over the entire dataset and the accuracy of the predicted labels (= given labels) over the consensus set for AG News Subset. Just like in Table 3, the accuracy of the predicted labels over the consensus set is significantly more than the accuracy of the predicted and given labels over the entire dataset. This explains why consensus-based RT performs the best, even though the consensus set is much smaller than the full dataset ($\sim 28\%$, $32\%$ and $38\%$ of the entire training set for the three values of $\epsilon$).
Table 11: AG News Subset. Accuracies of predicted labels and given labels over the entire dataset and accuracies of predicted labels (= given labels) over the consensus set. Conclusions are the same as in Table 3.
0.3 $53.20\pm 2.82$ $32.52\pm 2.05$ $\bm{61.81}\pm 2.66$
0.5 $66.78\pm 1.31$ $35.5\pm 0.14$ $\bm{76.48}\pm 0.93$
0.8 $79.98\pm 0.80$ $42.53\pm 0.13$ $\bm{89.59}\pm 0.43$
Appendix J Consensus-Based Retraining Does Better than Confidence-Based Retraining
Here we compare full and consensus-based RT against another strategy for retraining which we call confidence-based retraining (RT). Specifically, we propose to retrain with the predicted labels of the samples with the top $50\%$ margin (i.e., highest predicted probability - second highest predicted probability); margin is a measure of the model’s confidence. This idea is similar to self-training’s method of sample selection in the semi-supervised setting (Amini et al., 2022). In Tables 12 and 13, we show results for CIFAR-10 and CIFAR-100 (in the same setting as Section 5 and Appendix H) with the smallest value of $\epsilon$ from Tables 1 and 2, respectively. Notice that consensus-based RT is clearly better than confidence-based RT.
@@ Appendix I Accuracy over the Entire Dataset and over the Consensus Set in the Case of AG News Subset
Appendix K Retraining is Beneficial Even Without a Validation Set
In all our previous experiments, we assumed access to a small clean validation set. Here we show the results of training ResNet-34 on CIFAR-100 when we do not have access to a validation set and train for 100 epochs, and also compare them against the corresponding results with a validation set where we trained for 40 epochs (Table 4 in Section 5). The point of these experiments is to show that retraining can offer gains even without a validation set, in which case we train for too many epochs and expect overfitting. Please see the results and discussion in Table 14.
Table 14: CIFAR-100 w/ ResNet-34 without (top) and with (bottom) a validation set. Test set accuracies (mean $\pm$ standard deviation). Note that retraining (in particular, consensus-based RT) leads to improvement even without a validation set. However, the performance and amount of improvement obtained with retraining are worse in the absence of a validation set due to a higher degree of overfitting. This is not surprising.
Without Val. Set $\epsilon$ Baseline Full RT Consensus-based RT
@@
With Val. Set (Same as Table 4) $\epsilon$ Baseline Full RT Consensus-based RT
4 $37.53\pm 1.58$ $39.33\pm 1.37$ $\textbf{43.87}\pm 1.62$
5 $51.13\pm 0.69$ $52.33\pm 0.38$ $\textbf{55.43}\pm 0.42$
Appendix L Beyond Label DP: Evaluating Retraining in the Presence of Human Annotation Errors
Even though our empirical focus in this paper has been label DP training, retraining (RT) can be employed for general problems with label noise. Here we evaluate RT in a setting with “real” label noise due to human annotation. Specifically, we focus on training a ResNet-18 model (without label DP to be clear) on the CIFAR-100N dataset introduced by Wei et al. (2021) and available on the TensorFlow website. CIFAR-100N is just CIFAR-100 labeled by humans; thus, it has real human annotation errors. The experimental setup and details are the same as CIFAR-100 (as stated in Section 5 and Appendix H); the only difference is that here we use initial learning rate $=0.01$.
In Table 15, we list the test accuracies of the baseline which is just standard training with the given labels, full RT and consensus-based RT, respectively. Even here with human annotation errors, consensus-based RT is beneficial.